# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

### v2

@@ 3.4. Physics-aware Diffusion Transformer
[…]
Figure 5. Qualitative comparison with other SOTA methods on the UIEBD and LSUI datasets.
Figure 6. More visual results from real-world scenarios.
Figure 7. Limitations. The output images refer to the reconstruction images of PPG branch.
4. Experiments
4.1. Experimental Settings
Implementation details. Our network, implemented using PyTorch 1.7, underwent training and testing on an NVIDIA GeForce RTX 3090 GPU. We employed the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was established at 0.0001. During the training stage, to balance the batch size and image size, the batch and image size are set to 2 and 256 × 256, respectively. The number of training iterations reached one million. The pixel values of the image are normalized to [-1,1]. The time step of the diffusion model is set to 2000. In the testing stage, following the setting in (Tang et al., 2023), the input size of the image is 256 × 256. By using the skip sampling strategy, and the sampling times are set to 10 times to balance the performance and runtime.
Datasets. We utilize the real-world UIEBD dataset (Peng et al., 2023) and the LSUI dataset (Li et al., 2020) for training and evaluating our model. The UIEBD dataset comprises 890 underwater images with corresponding labels. Out of these, 700 images are allocated for training, and the remaining 190 are designated for testing. LSUI dataset contains 5004 underwater images and their corresponding high-quality images. Compared with UIEBD, LSUI contains diverse underwater scenes, object categories and deep-sea and cave images. In the paper, The LSUI dataset is randomly partitioned into 4500 images for training and 504 images for testing. In addition, to verify the generalization of PA-Diff, we use non-reference benchmarks U45 (Li et al., 2019), which contains 45 underwater images for testing.
Comparison methods. We conduct a comparative analysis between PA-Diff and seven state-of-the-art (SOTA) UIE methods, namely UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UWCNN (Anwar et al., 2018), UIEWD (Ma and Oh, 2022), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023). To ensure a fair and rigorous comparison, we utilize the provided source codes from the respective authors and adhere strictly to the identical experimental settings across all evaluations.
Evaluation Metrics. We employ standard full-reference image quality assessment metrics, including PSNR and SSIM (Wang et al., 2004), for quantitative comparisons at both pixel and structural levels. Higher PSNR and SSIM values indicate superior image quality. Additionally, LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) are utilized to evaluate perceptual performance. Lower LPIPS and FID scores signify a more effective UIE approach. For non-reference benchmarks U45, we introduce UIQM (Panetta et al., 2015) and UCIQE (Yang and Sowmya, 2015) to evaluate our method.
Table 2. Ablation study on LSUI dataset. $T^{c}$ and $B^{c}$ refer to transmission map and background light prior information in PPG branch, respectively. GBB means Gaussian blur module. INR represents our INR branch. PSE refers to the periodic spatial encoding (Lee and Jin, 2022), aiming to better recover high-frequency details. CAM and PPU means our designed cross attention module and physics perception unit , respectively. $GM-FFN$ means multi-scale operation in the feed-forward network.
[…]
A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 28.35 0.8893
B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 27.71 0.8974
C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.92 0.9071
D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.24 0.9112
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 24.50 0.9186
F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 24.92 0.9176
G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 23.66 0.9241
H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 22.85 0.9374
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 22.13 0.9378
4.2. Results and Comparisons
Table 1 shows the quantitative results compared with different SOTA methods on the UIEBD, LSUI and U45 datasets, including with UIECˆ2-Net (Wang et al., 2021), Water-Net (Peng et al., 2023), UIEWD (Ma and Oh, 2022), UWCNN (Anwar et al., 2018), U-color (Li et al., 2021), U-shape (Li et al., 2020), and DM-water (Tang et al., 2023). We mainly use PSNR, SSIM, LPIPS and FID as our quantitative indices for UIEBD and LSUI datasets, and UIQM and UCIQE for non-reference dataset U45. The results in Table 1 show that our algorithm outperforms state-of-the art methods obviously, especially in terms of perceptual metrics. To better validate the superiority of our methods, Fig. 5 shows the visual comparison results with other methods on the UIEBD and LSUI datasets. The some examples are randomly selected on the UIEBD and LSUI datasets. Furthermore, in order to more effectively validate the generalization performance of our method on real-world images, we present enhancement results from real scenes in Figure 6. Our methods consistently generate natural and better visual results, strongly proving that PA-Diff has good generalization performance for real-world underwater images.
4.3. Ablation Study
@@
To evaluate the impact of each strategy on the diffusion model, we conduct a comprehensive ablation study. Table 2 shows the ablation results on the LSUI dataset. We use a full-reference assessment metric (SSIM) and perceptual assessment metric (FID) to evaluate the ablation study, respectively. Model A means that we just employ the most basic diffusion transformer structure. Firstly, we investigate the influence of the PPG branch on the diffusion process. Compared with model A, model B, C and D achieve better results, suggesting that the prior knowledge of physics can help diffusion model to better restore image. Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T^{c}$ plays a highly effective role in guiding the diffusion process. Model D outperforms C, suggesting that GBB is also useful to exploit more helpful the background light information for the diffusion process. Subsequently, we explore the impact of the INR branch on the performance of the diffusion model. Model E outperforms D, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models. Model F outperforms E, implying that the periodic spatial encoding can enhance the representation ability of INR branch. Subsequently, for models G and H, we respectively delve into the effects of our designed CAM and PPU on the model performance. Both models G and H yield improved outcomes, proving the significant role that CAM and PPU play in extracting more valuable feature-level physical information. Finally, Ours achieves the best performance, proving that our all contribution is effective for UIE tasks. Moreover, we also examine the applicability of PPU in other baseline models. Table 3 showcases the adaptability of PPU across different UIE methods, demonstrating its generalizability potential to exploit the beneficial feature-level physics information.
Limitations. As illustrated in Figure 7, there remains a discernible discrepancy between the reconstructed images from the PPG branch and the corresponding input images, indicating an existing gap between the physical information generated by our PPG branch and the ideal or ground-truth physical representation. Consequently, how to generate more accurate a prior knowledge of physics remains a problem we need to research in the future.
Table 3. Ablation study with the PPU on UIEBD.
@@
w/o PPU(PSNR/SSIM) 17.41/0.7769 21.88/0.8194 21.25/0.8453
w PPU(PSNR/SSIM) 19.64/0.8157 21.23/0.8331 21.89/0.8570
5. Conclusion
In this paper, we develop a novel UIE framework, namely PA-Diff. With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. To the best of our knowledge, it is the first the diffusion model based on physical perception in image enhancement tasks. Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into any deep learning network. PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments can prove that each of our contributions is effective.

### v3

@@ IV-B Results and Comparisons
A $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.14 0.852
B $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.67 0.857
C $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.73 0.865
D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 25.94 0.868
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 26.47 0.873
F $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ 26.51 0.872
G $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 26.79 0.880
H $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ 27.04 0.883
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 27.25 0.889
TABLE VI: Comparison results of inference speed and Params.
Method WaterNet U-color Ushape DM-water PA-Diff
Inference time(s) 0.61 2.75 0.03 0.14 0.23
Params(M) 24.81 157.4 65.60 10.69 41.76
TABLE VII: Params analysis of each branch in PA-Diff.
Method PPG branch INR branch PDT branch PA-Diff (all)
Params(M) 0.0302 0.1487 41.31 41.76
IV-C Ablation Study
Effects of Diffusion Technology. Firstly, to validate the value of the diffusion technique within our framework, Table IV presents the ablation study on the application of diffusion technology. The proposed model with diffusion technique achieves better results, demonstrating the significant performance improvements it brings to the model.
Analysis of Model Complexity. Table VI shows a comparison of inference speeds and model parameters with the previous methods, indicating that our method demonstrates competitive model parameters. Table VII presents the network parameters for each branch, clearly indicating that the PPG and INR branches, as auxiliary networks for prior generation, have a minimal number of parameters and exhibit low complexity. However, we employ a transformer-based denoising network in the PDT branch, resulting in relatively large parameters. Therefore, our future endeavors will concentrate on investigating lighter-weight denoising networks.
@@ IV-C Ablation Study
Impact of Different Framework. To evaluate the impact of each strategy on the diffusion model, we conduct a comprehensive ablation study. Table V shows the ablation results on the LSUI dataset, and Figure 5 presents the visual results of the prior knowledge. Model A means that we just employ the most basic diffusion transformer structure. Firstly, we investigate the influence of the PPG branch on the diffusion process. Compared with model A, model B, C and D achieve better results, suggesting that physical prior knowledge helps the diffusion model better restore underwater images. Particularly, Model B exhibits a remarkable advancement in terms of PSNR metric, suggesting that the transmission map $T_{c}$ plays a highly effective role in guiding the diffusion process. Subsequently, we explore the impact of the INR branch on the performance of the diffusion model. Model E outperforms D, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models. Model F outperforms E, implying that the periodic spatial encoding can enhance the representation ability of INR branch. Finally, Ours achieves the best performance, proving that all our proposed components are effective for UIE tasks.
Analysis and Discussion. We conduct an exhaustive analysis of the individual contributions of each core component to the holistic performance by visually presenting the results of ablation studies, as illustrated in Figure 9. Here, Model b represents the employment of the fundamental diffusion transformer architecture alone, serving as our baseline. Compared to the full model, the results of Model b show that the visual effects obtained by using only the basic diffusion model can be unrealistic. Models c and d are designed to investigate the influence of incorporating prior physical knowledge—embodied by the transmission map $T_{c}$ and background light information $B_{c}$, respectively—into the full model framework. Compared to model d, the results for model c are worse, mainly in terms of poorer recovery in some areas and less good colour recovery, suggesting that the transmission map $T_{c}$ we introduced plays a very important role. Model e, on the other hand, focuses on elucidating the impact of integrating an Implicit Neural Representation (INR) branch. Compared to the full model, the results of Model e show that the visual effects obtained by using only the basic diffusion model can be bad, suggesting that INR branch can improve the performance and reduce the difficulty of restoration for diffusion models. Lastly, Model F is pivotal in examining the role of the PPU module within the full model context. The visual evidence gathered unmistakably highlights the crucial role each segment of our proposed model plays in achieving optimal performance.
Impact of Physical Priors. As illustrated in Figure 10, we show a comparative analysis of diffusion-based underwater image restoration with and without physical priors. The results demonstrate that, in the absence of physical priors, the diffusion model exhibits noticeable deviations in both global color correction and local detail reconstruction, particularly under severely degraded underwater conditions. In contrast, incorporating physical priors provides a robust guidance signal throughout the diffusion process, effectively alleviating error accumulation during reverse denoising. This design enables the diffusion model to better align with the underlying degradation mechanism, thereby improving its generalization capability and restoration robustness.
The Plug-and-play Capability of PPU. Furthermore, we explore the application of the PPU in other models. Table VIII showcases the adaptability of PPU across different methods, demonstrating its universal potential across diverse methods to exploit the beneficial feature-level physics information.
Impact of Different PPU Structure. We conduct an in-depth analysis of the key design choices in the proposed PPU. Specifically, we examine two critical physical parameters, $\tilde{\boldsymbol{t}_{1}}$ and $\tilde{\boldsymbol{t}_{2}}$, which are closely related to the transmission map. As shown in Table IX, removing either parameter leads to a significant performance degradation, validating the effectiveness of the current PPU design. Furthermore, we include comparisons with alternative feature-level physics modeling modules, such as the PDU used in C2PNet [70]. The results demonstrate that our PPU consistently achieves superior performance, indicating that the proposed PPU is more suitable for underwater image enhancement.
Choice of Attention in the PPG Branch. We perform ablation studies on the specific attention mechanisms adopted in the PPG branch, as reported in Table X. Variants that remove either channel attention or spatial attention are evaluated. The results show a consistent, albeit slight, performance drop when either component is removed, supporting the effectiveness of our attention design.
A Lightweight Variant of Our PA-Diff. we design a lightweight variant of our model by reducing the channel width and the number of modules in the transformer-based PDT. As reported in Table XI, the proposed lightweight PA-Diff (Mini) achieves performance comparable to the full model across all evaluation metrics, while using only about one quarter of the parameters of the full model. Moreover, when compared with DM-Water under a similar parameter budget, our lightweight model consistently delivers superior results, further validating the effectiveness and scalability of the proposed framework.
Extention on Downstream Tasks. As shown in Figure 11, we provide a multi-view visual comparison using the UwMVS real-world test set [50], which further validate the effectiveness of our proposed method in improving reconstruction quality.
TABLE VIII: Ablation study with the PPU on UIEBD.
@@
w/o PPU(PSNR/SSIM) 17.41/0.7769 22.78/0.894 18.15/0.833
w PPU(PSNR/SSIM) 19.64/0.8157 22.93/0.897 20.54/0.848
TABLE IX: Ablation with the choice of PPU structure on the LSUI dataset.
LSUI
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
w/o $\tilde{\boldsymbol{t}_{1}}$ 26.31 0.880 0.130 24.61 0.590 4.306
w/o $\tilde{\boldsymbol{t}_{2}}$ 25.88 0.876 0.135 33.96 0.585 4.232
PDU [70] 26.94 0.878 0.126 22.76 0.593 4.318
PA-Diff 27.25 0.889 0.103 20.46 0.594 4.393
TABLE X: Ablation with the attention mechanisms of the PPG Branch on the LSUI dataset. CA and SA denote channel attention and spatial attention, respectively.
LSUI
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ UCIQE $\uparrow$ UIQM $\uparrow$
w/o CA 27.11 0.885 0.110 22.14 0.592 4.372
w/o SA 27.08 0.879 0.104 21.83 0.589 4.381
PA-Diff 27.25 0.889 0.103 20.46 0.594 4.393
TABLE XI: Comparison with lightweight variants on the LSUI dataset.
LSUI
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ UIQM $\uparrow$ Params(M) $\downarrow$
Ushape 23.66 0.845 0.158 4.351 65.60
DM-water 26.95 0.881 0.143 4.257 10.69
PA-Diff (Mini) 27.18 0.889 0.123 4.378 11.58
PA-Diff (Ours) 27.25 0.889 0.103 4.393 41.76
Limitations. As illustrated in Figure 12, there remains a discernible discrepancy between the reconstructed images from the PPG branch and the corresponding input images, indicating an existing gap between the physical information generated by our PPG branch and the ideal or ground-truth physical representation. Consequently, how to generate more accurate physical prior knowledge remains a challenge for future research. Regarding the extreme low-light conditions, we add a new set of qualitative results on the LSUI dataset. As shown in Figure 13, while our method is able to achieve noticeable overall enhancement in terms of brightness and color consistency, the recovery of fine-grained details remains challenging due to severe information loss. Moreover, the performance under such extreme conditions is further constrained by the limited quality of the available data, as the ground-truth images themselves are often noisy or poorly illuminated. Thus, we also plan to explore self-supervised learning paradigms and the construction of higher-quality datasets to better address extreme low-light underwater scenarios.
V Conclusion
In this paper, we develop a novel UIE framework, namely PA-Diff. With utilizing physics prior information to guide the diffusion process, PA-Diff can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. To the best of our knowledge, it is the first diffusion model based on physical perception for UIE tasks. Our designed PPU is a plug-and-play universal module to exploit the beneficial feature-level physics information, which can be integrated into various deep learning network. PA-Diff shows SOTA performance on UIE task, and extensive ablation experiments validate the effectiveness of each proposed component.